OpenStudy (anonymous):

The perimeter of the rectangle 26ft. The length is 5ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length y be the width. 2x+2y=_ x=y+__ What is the Length? __ft What is the Width? __Ft

OpenStudy (shamil98):

You have all the givens to solve. p = 26 l = 5 + w 26= 2l + 2w 26 = 2(5+w) + 2w 26 = 10 + 2w + 2w 26 = 10 +4w 16 = 4w w = 4 now just plug 4 into the equation for L , l = 5 + 4 l = 9